Assuming that the chemical reaction obeys zero-order reaction, what is the time when of 14 M of substance A will decrease to 2 M, and the chemical reaction has k = 4.52x10 2 M/s?

3A → 2 B + C
(zero-order kinetics)
1
Expert's answer
2020-10-19T14:24:01-0400

Given CA0 = 14 M = initial concentration of the reactant at time t=0

CA = 2M= final Concentration of the reactant at the time t

Rate constant K = 4.52×102M/S

Time=t=?

Chemical reaction obeys zero order reaction

i.e n=0 (zero order kinetics)

Order of reaction with respect to A is zero i.e n= 0

3A------->2B + C

Rate equation

-rA = KCAn

Put n=0

-dCA/dt = KCA0

-dCA/dt = K

So the Integrated form of above equation is

CA0 - CA = Kt

t = (CAo - CA)/K

t =(14M - 2M)/(4.52×102M/S)

t= 0.0265 sec